WHEN THREE BLOCKS OF(DIFFERENT MASSES) ARE PLACED ON TOP OF ONE ANOTHER SUCH THAT THERE IS FRICTION(FRICTION COEFFICIENT DIFFERENT)BETWEEN ALL BLOCKS AND ALSO THE GROUND THEN HOW SHOULD I APPROACH THE PROBLEM WHEN THERE IS A FORCE APPLIED ON THE 2ND BLOCK AND I HAVE FIND THE ACCELERATIONS & THE FRICTIONAL FORCES?


WHEN I USE FRICTION=kN ;


I GET THE WRONG ANSWERS. PLEASE HELP!

Grade:

1 Answers

Rathod Shankar AskiitiansExpert-IITB
13 years ago

1) consider the block which is under force. so apply newtons method for force equation.

 

sa applied force is F on that block.

here u= co efficeint

N= normal force

friction force f= U*N

F-f1-f3=m2*a

F-u1 * m1 *g-u3*m3*g= m2*a 2  ...... u1 is between m2 amd m1....u3 is between m2 and m3

 

apply same equations on 1st and 3rd blocks. the friction force direction is opposite to f1 and f3.

f3=m3*a3 = u3* m3*g

 

f3=m1*a1 = u1* (m3+m2+m1)*g

 

 

thanks

Rathod Shankar

IIT Bombay